OpenStudy (anonymous):

I FAN AND MEDAL Solve the following system of equations: x - 2y = 14 x + 3y = 9 A (1, 12) B (-1, -12) C (12, -1) D (12, 1)

OpenStudy (anonymous):

I also need help with this one :) but no pressure it's okay if no one can answer it Determine which system below will produce infinitely many solutions. A) 2x + 5y = 24 2x + 5y = 42 B) 3x - 2y = 15 6x + 5y = 11 C) 4x - 3y = 9 -8x + 6y = -18 D) 5x - 3y = 16 -2x + 3y = -7

OpenStudy (anonymous):

ill do the first stew you do the second one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

sorry i was in chat let me pay attention

OpenStudy (anonymous):

Oh its okay

OpenStudy (anonymous):

I'm not sure of where to start for these problems

OpenStudy (anonymous):

sorry i have the answer but i dont want to give you it right away i got banned for doing that the last time

OpenStudy (anonymous):

Oh it's okay! & the first

OpenStudy (anonymous):

What do you mean by stew lol

OpenStudy (anonymous):

helloo @My_Yich

OpenStudy (anonymous):

Ok, so you have to isolate a variable so you can start to solve it. Since the x value is already in a 1:1 ratio with itself, lets start with that one. Multiply the bottom by -1

OpenStudy (anonymous):

so now you have x - 2y = 14 -x - 3y = -9

OpenStudy (anonymous):

then what you want to do is Add the two equations together: Add the two equations together: -5y = 5 y = -1

OpenStudy (anonymous):

@Stewpeas_0831

OpenStudy (anonymous):

Now you can plug this back in to the first equation to find x. x - 2(-1) = 14 x + 2 = 14 x = 12

OpenStudy (anonymous):

To check to see if this is correct, plug them both into the second equation to make sure it makes sense. (12) + 3(-1) = 9

OpenStudy (anonymous):

soo whats the answer

OpenStudy (anonymous):

(12,-1)! Thanks so much that helped a lot!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

multiply the first equation by 2 4x - 3y = 9 -->(2) -8x + 6y = -18 8x - 6y = 18 (result of multiplying by 2) -8x + 6y = -18 add both equations together and then you will get your answer